# Anisotropic Collective Excitations of Bose Gases in Modified Newtonian Dynamics

### v1

@@ Abstract
Collective excitations are fundamental in quantum many-body physics, yet their spectra have traditionally been studied within Newtonian dynamics. In this Letter, we investigate collective excitations in Bose gases under Modified Newtonian Dynamics (MOND). We derive an anisotropic excitation spectrum in the MOND regime. This anisotropy arises directly from the intrinsic nonlinear structure of the MOND Poisson equation, forming a distinctive signature of the modified gravitational response. We then analyze the Jeans instability, obtaining analytic expressions for the direction-dependent critical wavelength and mass. These results advance our understanding of collective behavior in quantum systems under modified dynamics and establish clear theoretical signatures for testing MOND-like effects in quantum simulators.
Introduction.
Collective excitations represent a fundamental concept in quantum many-body physics. They characterize the intrinsic quantum nature of many-particle systems and provide key insights into emergent phenomena such as superfluidity and superconductivity Griffin (1993); Bruus and Flensberg (2004); Bloch et al. (2008); Coleman (2015). These low-energy modes describe how a quantum system responds to small perturbations. Their properties have traditionally been studied within the framework of Newtonian dynamics. In this work, we extend this paradigm by investigating collective excitations in a Bose gas under Modified Newtonian Dynamics (MOND). This exploration is motivated by the possibility of uncovering novel physical behavior in quantum many-body systems that fundamentally distinguishes MOND from its Newtonian counterpart.
@@ Introduction.
Direct tests of MOND in terrestrial laboratories are challenging due to the extremely low acceleration scale $a_{0}$. However, quantum simulators, particularly ultracold atomic Bose-Einstein condensates (BECs) Dalfovo et al. (1999); Fedichev and Fischer (2003, 2004); Boehmer and Harko (2007); Pethick and Smith (2008); Lewenstein et al. (2012); Steinhauer (2014); Das and Bhaduri (2015); Steinhauer (2016); Gross and Bloch (2017); Chä and Fischer (2017); Eckel et al. (2018); Barceló (2019); Muñoz de Nova et al. (2019); Kolobov et al. (2021); Holanda Ribeiro and Fischer (2023); Pal and Fischer (2024); Haddad (2025), offer a promising platform to emulate certain features of MOND-inspired physics in a controlled setting. By engineering appropriate long-range interactions O’dell et al. (2000); Giovanazzi et al. (2001); Choi (2002); Chalony et al. (2013); Defenu et al. (2023), these systems may mimic the low-acceleration behavior characteristic of MOND. Self-gravitating BECs have been studied as analog models for dark matter halos. Their collective excitations in the Newtonian limit are described by the dispersion relation Chavanis (2011, 2017, 2020, 2025)
@@
where the terms correspond to quantum pressure, short-range contact interactions, and gravitational attraction, respectively. Here, $n_{0}$ is the condensate density, $\rho_{0}=mn_{0}$ the mass density, $g$ the interaction strength, and $G$ Newton’s constant.
Recent theoretical work has begun exploring ground-state properties of BECs under MOND-like logarithmic potentials, revealing distinctive features such as significant cloud enlargement compared to Newtonian traps Liu (2026). These findings motivate a deeper investigation into dynamical properties, particularly the full collective excitation spectrum, within a consistent MOND framework.
@@
Model.
The system consists of a self-gravitating BEC described by a macroscopic wave function $\psi(\mathbf{r},t)$ and a gravitational potential $\Phi(\mathbf{r},t)$. Its dynamics follows from the Lagrangian density
@@
Bogoliubov-de Gennes Theory.
To study small-amplitude collective modes, we expand the wave function and gravitational potential around their equilibrium values:
@@
The full anisotropic landscape is displayed in Fig. 2 as a contour plot of $\tilde{\omega}^{2}$ in the $\tilde{k}-$$\theta$ plane. The red dashed contour marks the stability boundary $\tilde{\omega}^{2}=0$, which separates the parameter space into two regions: the stable collective‑excitation region on the right ($\tilde{\omega}^{2}>0$) and the unstable, collapse‑prone region on the left ($\tilde{\omega}^{2}<0$). The shape of this boundary clearly shows that the critical wavenumber for stability decreases as $\theta$ increases from $\theta=0$ to $\theta=\pi/2$, confirming the directional dependence of the Jeans instability.
@@
Jeans Instability Analysis.
In the long‑wavelength limit $\tilde{k}\to 0$, Eq. (25) gives
@@
which likewise exhibits a clear $\theta$‑dependence. Hence, the anisotropy of the Jeans scale is generic consequence of the MOND nonlinearity, persisting in both quantum and classical regimes.
@@
The angular dependence of the Jeans mass has important implications for structure formation in self‑gravitating systems under MOND. Unlike Newtonian gravity, where collapse occurs isotropically, MOND predicts that gravitational instability is more efficient in directions perpendicular to the background field, favoring the formation of anisotropic structures. On the other hand, the anisotropic instability can be a hallmark of MOND and provides a distinctive signature that could be tested in quantum simulators with engineered long‑range interactions.
Summary and outlook.
We have derived the collective excitation spectrum of a self‑gravitating Bose‑Einstein condensate in the MOND framework. The dispersion relation exhibits a clear anisotropy, depending on the angle between the perturbation wavevector and the background gravitational field. This anisotropy stems from the nonlinearity of the MOND Poisson equation and disappears in the Newtonian limit. In the deep‑MOND regime, the Jeans instability becomes direction‑dependent, with perpendicular perturbations being more unstable than parallel ones.
Our work extends the study of quantum fluids to the low‑acceleration regime described by MOND. The anisotropic excitation spectrum offers a distinct signature that could distinguish MOND‑like effects from standard Newtonian behavior. Although derived here for a BEC, the same anisotropy should appear in other quantum liquids and in classical systems, because its origin lies in the modified gravitational response. This universality makes the effect a promising target for experimental simulation.

### v2

@@ Abstract
Collective excitations are fundamental in quantum many-body physics, yet their spectra have traditionally been studied within Newtonian dynamics. In this paper, we investigate collective excitations in Bose gases under Modified Newtonian Dynamics (MOND). We derive an anisotropic excitation spectrum in the MOND regime. This anisotropy arises directly from the intrinsic nonlinear structure of the MOND Poisson equation, forming a distinctive signature of the modified gravitational response. We then analyze the Jeans instability, obtaining analytic expressions for the direction-dependent critical wavelength and mass. These results advance our understanding of collective behavior in quantum systems under modified dynamics and establish clear theoretical signatures for testing MOND-like effects in quantum simulators.
Introduction. Collective excitations represent a fundamental concept in quantum many-body physics. They characterize the intrinsic quantum nature of many-particle systems and provide key insights into emergent phenomena such as superfluidity and superconductivity Griffin (1993); Bruus and Flensberg (2004); Bloch et al. (2008); Coleman (2015). These low-energy modes describe how a quantum system responds to small perturbations. Their properties have traditionally been studied within the framework of Newtonian dynamics. In this work, we extend this paradigm by investigating collective excitations in a Bose gas under Modified Newtonian Dynamics (MOND). This exploration is motivated by the possibility of uncovering novel physical behavior in quantum many-body systems that fundamentally distinguishes MOND from its Newtonian counterpart.
@@
Direct tests of MOND in terrestrial laboratories are challenging due to the extremely low acceleration scale $a_{0}$. However, quantum simulators, particularly ultracold atomic Bose-Einstein condensates (BECs), offer a versatile platform to emulate features of gravitational physics in a controlled setting Dalfovo et al. (1999); Fedichev and Fischer (2003, 2004); Boehmer and Harko (2007); Pethick and Smith (2008); Lewenstein et al. (2012); Steinhauer (2014); Das and Bhaduri (2015); Steinhauer (2016); Gross and Bloch (2017); Chä and Fischer (2017); Eckel et al. (2018); Barceló (2019); Muñoz de Nova et al. (2019); Kolobov et al. (2021); Holanda Ribeiro and Fischer (2023); Pal and Fischer (2024); Haddad (2025). A concrete proposal for mimicking a Newtonian potential was put forward by O’Dell et al. O’dell et al. (2000), who showed that six appropriately arranged off-resonant laser beams can induce an attractive $1/r$ interaction between neutral atoms. In the near-zone limit, the faster-decaying $1/r^{3}$ dipole-dipole terms average out, leaving the desired long-range $1/r$ potential. Although a self-bound condensate with such an interaction has not yet been realized experimentally, the required physical parameters have been analyzed in detail Giovanazzi et al. (2001), and the theoretical properties of the resulting system have been studied extensively Cartarius et al. (2008). Self-gravitating BECs themselves have been widely investigated as analog models for dark-matter halos Chavanis (2011, 2017, 2020, 2025). Their collective excitations in the Newtonian limit are described by the dispersion relation
@@
where the terms correspond to quantum pressure, short-range contact interactions, and gravitational attraction, respectively. Here, $n_{0}$ is the condensate density, $\rho_{0}=mn_{0}$ the mass density, $g$ the interaction strength, and $G$ Newton’s constant. Recent theoretical work has begun exploring ground-state properties of BECs under MOND-like logarithmic potentials, revealing distinctive scaling laws and significant cloud enlargement compared to Newtonian traps Liu (2026). These findings motivate a deeper investigation into dynamical properties, particularly the full collective excitation spectrum, within a consistent MOND framework.
@@
Model. The system consists of a self-gravitating BEC described by a macroscopic wave function $\psi(\mathbf{r},t)$ and a gravitational potential $\Phi(\mathbf{r},t)$. Its dynamics follows from the Lagrangian density
@@
Bogoliubov-de Gennes Theory. To study small-amplitude collective modes, we expand the wave function and gravitational potential around their equilibrium values:
@@
The full anisotropic landscape is displayed in Fig. 2 as a contour plot of $\tilde{\omega}^{2}$ in the $\tilde{k}-$$\theta$ plane. The red dashed contour marks the stability boundary $\tilde{\omega}^{2}=0$, which separates the parameter space into two regions: the stable collective-excitation region on the right ($\tilde{\omega}^{2}>0$) and the unstable, collapse-prone region on the left ($\tilde{\omega}^{2}<0$). The shape of this boundary clearly shows that the critical wavenumber for stability decreases as $\theta$ increases from $\theta=0$ to $\theta=\pi/2$, confirming the directional dependence of the Jeans instability.
@@
Jeans Instability Analysis. In the long-wavelength limit $\tilde{k}\to 0$, Eq. (25) gives
@@
which likewise exhibits a clear $\theta$-dependence. Hence, the anisotropy of the Jeans scale is generic consequence of the MOND nonlinearity, persisting in both quantum and classical regimes.
@@
The angular dependence of the Jeans mass has important implications for structure formation in self-gravitating systems under MOND. Unlike Newtonian gravity, where collapse occurs isotropically, MOND predicts that gravitational instability is more efficient in directions perpendicular to the background field, favoring the formation of anisotropic structures. On the other hand, the anisotropic instability can be a hallmark of MOND and provides a distinctive signature that could be tested in quantum simulators with engineered long-range interactions.
Summary and outlook. We have derived the collective excitation spectrum of a self-gravitating Bose-Einstein condensate in the MOND framework. The dispersion relation exhibits a clear anisotropy, depending on the angle between the perturbation wavevector and the background gravitational field. This anisotropy stems from the nonlinearity of the MOND Poisson equation and disappears in the Newtonian limit. In the deep-MOND regime, the Jeans instability becomes direction-dependent, with perpendicular perturbations being more unstable than parallel ones.
Our work extends the study of quantum fluids to the low-acceleration regime described by MOND. The anisotropic excitation spectrum offers a distinct signature that could distinguish MOND-like effects from standard Newtonian behavior. Although derived here for a BEC, the same anisotropy should appear in other quantum liquids and in classical systems, because its origin lies in the modified gravitational response. This universality makes the effect a promising target for experimental simulation.